# A single-layer framework of variational tensor network states

### v1

@@ Abstract
We propose a single-layer tensor network framework for the variational determination of ground states in two-dimensional quantum lattice models. By combining the nested tensor network method [Phys. Rev. B 96, 045128 (2017)] with the automatic differentiation technique, our approach can reduce the computational cost by three orders of magnitude in bond dimension, and therefore enables highly efficient variational ground-state calculations. We demonstrate the capability of this framework through two quantum spin models: the antiferromagnetic Heisenberg model on a square lattice and the frustrated Shastry-Sutherland model. Even without GPU acceleration or symmetry implimention, we have achieved the bond dimension of 9 and obtained accurate ground-state energy and consistent order parameters compared to prior studies. In particular, we confirm the existence of an intermediate empty-plaquette valence bond solid ground state in the Shastry-Sutherland model. We have further discussed the convergence of the algorithm and its potential improvements. Our work provides a promising route for large-scale tensor network calculations of two-dimensional quantum systems.
@@ I Introduction
Two-dimensional (2D) strongly correlated systems, such as interacting fermionic systems and frustrated quantum magnets, pose significant challenges for theoretical studies. Consequently, developing efficient and accurate numerical methods for these models constitutes a crucial branch of theoretical condensed matter physics. In recent decades, tensor network methods XiangBook ; SimonBook have attracted increasing attention. By combining the tensor-network ansatz PEPS2004 ; OrusNatRev stemming from the quantum entanglement, the contraction strategy originating from renormalization group Orus2014 ; RanBook , and modern optimization techniques developed in artificial intelligence OptBook , tensor network methods have been successfully applied to diverse phenomena, including quantum magnetism KagomePRL2017 ; J1J2-LWY ; NingXi-SSM2023 ; LY-PRB2025 , superconductivity TNStJ ; TNSHubbard ; Corboz-MulCTMRG , topological order TNStopo ; CSL , and quantum field theory CMPS ; CTNS .
For studying ground states of 2D quantum systems using tensor networks, three strategies are commonly adopted. The first maps the 2D quantum model to a three-dimensional classical one QCmap , converting the quantum expectation value calculation into a statistical physics problem HOTRG . The second employs imaginary-time evolution with wave function update schemes SU ; CU ; FU ; FFU to obtain the ground state $|\Psi_{g}\rangle$ in the tensor-network representation, and then evaluate the expectation value $\langle\hat{O}\rangle$ of the operator $\hat{O}$ by contracting the related tensor networks. The Third one is the straightforward variational minimization: Starting from an arbitrary tensor network state $|\Psi\rangle$, the ground state of the system $|\Psi_{g}\rangle$ and its energy $E_{g}$ are obtained by minimizing the energy estimator
@@
with respect to tensor parameters PEPS2004 . Recently, this approach has become increasingly popular due to the advent of the automatic differentiation technique ADreview , which, as a mature and widely applied algorithm NeuralBook , can efficiently calculate the gradient $\frac{dE}{d|\Psi\rangle}$ LW-PTNS2019 and easily integrate with gradient-based optimization method OptBook via various open-source packages Pytorch ; Zygote .
@@
A central challenge in all three approaches is the contraction of tensor networks. Unfortunately, contracting a general tensor network in more than one dimension is a #P\#P-complete hard problem PEPSComplex that cannot be performed exactly for an infinite system even with translational invariance. To this end, a series of approximation algorithms based on the renormalization group idea have been developed RanBook ; XiangBook , among which the corner transfer-matrix renormalization group (CTMRG) Nishino-CTMRG ; Vidal-CTM ; Corboz-MulCTMRG ; Vari-CTMRG has been widely used. In two dimensions, the cost of the CTMRG algorithm scales as $D^{10}$ in memory and $D^{12}$ in computation with the bond dimension $D$ when no symmetry or partial matrix decomposition is employed. The cost is extremely high for variational calculations, which typically require hundreds of rounds of such contractions. The problem becomes even more serious for minimization with the automatic differentiation since it needs huge memory to store intermediate data for the gradient ($\frac{dE}{d|\Psi\rangle}$) calculation. This issue pose severe limitation for the bond dimension $D$ in variational minimization employing automatic differentiation.
@@
To overcome this limitation, here we propose a single-layer framework of variational tensor network state calculation that incorporates the nested tensor network (NTN) method NTN2017 . By reformulating the double-layer structures of numerator and denominator in Eq. (1) into nested single-layer structures, the NTN method reduces the memory and computation cost of tensor network contraction by two and three orders of $D$, respectively. This allows us to perform variational minimization with significantly larger bond dimension. Our proposal works as follows: For any given state $|\Psi\rangle$, we utilize the NTN to calculate the energy defined in Eq. (1), then calculate the gradient $\frac{dE}{d|\Psi\rangle}$, and finally update the wave function $|\Psi\rangle$ using the gradient information as usual LW-PTNS2019 . This procedure can be iterated many times until the calculated energy converges.
@@
We validate this method on two benchmark models defined on infinite 2D lattices: the $S=1/2$ square-lattice antiferromagnetic Heisenberg model and the frustrated Shastry-Sutherland model. The former has an antiferromagnetically long-range ordered ground state HM-MC , while the latter hosts a series of ground states when tuning the strength of spin frustration $J/J^{\prime}$ Corboz-SSM ; NingXi-SSM2023 . In both cases, with the help of NTN, we can efficiently reach $D=9$ even without any GPU utilization or symmetry implementation, which is difficult in previous variational calculations LW-PTNS2019 ; NingXi-SSM2023 ; LY-PRB2025 . The obtained ground-state energy and order parameters are consistent with previous calculations, but with better convergence. In particular, we provide clear evidence for an empty-plaquette valence bond solid (VBS) phase in the ground-state phase diagram of the Shastry-Sutherland model.
@@
The rest of the paper is organized as follows. In Sec. II we first briefly review of the automatic differentiation-assisted variational calculation of tensor network state and the NTN method, then introduce our single-layer framework. In Sec. III, we present the numerical results of the Heisenberg model and the Shastry-Sutherland model in detail. Finally, we summarize and conclude in Sec. IV.
@@ II Method
In this section, we first review the basics of automatic differentiation-assisted variational calculation of tensor network representation of the ground state LW-PTNS2019 , the underlying double-layer structure in Eq. (1), and the NTN method NTN2017 , then we introduce the single-layer framework by combining the two parts together.
@@ II.1 Variational calculation assisted by automatic differentiation
In this work, we follow the variational minimization approach outlined in Sec. I. Starting from an initial $|\Psi(X_{0})\rangle$, we can use well-established techniques, such as the CTMRG Nishino-CTMRG ; Vidal-CTM ; Corboz-MulCTMRG ; Vari-CTMRG , to compute the energy $E(X)$. In this procedure, when the automatic differentiation technique is used. Then, as long as $E$ is obtained successfully, we can calculate the gradient $\frac{\partial{E(X)}}{\partial{X}}\big|_{X=X_{0}}$ conveniently by simply calling some open source packages, such as PyTorch Pytorch and Zygote Zygote , without any extra operation. After the gradient information is obtained, we can update $X$ using classic optimization algorithms, such as L-BFGS OptBook , to get a better trial $X_{1}$ and the corresponding wave function $|\Psi(X_{1})\rangle$. Similary, we can evaluate $E(X_{1})$, calculate $\frac{\partial{E(X)}}{\partial{X}}\big|_{X=X_{1}}$, and then obtain a better $X_{2}$ and $|\Psi(X_{2})\rangle$. This iteration can be repeated as many as we want until the gradient $\frac{\partial{E(X)}}{\partial{X}}$ is sufficiently small or the energy $E$ converges within some tolerance. The converged $X^{*}$ yields an approximate solution of Eq. (4) and thereby provides an approximate representation of $|\Psi_{g}\rangle$.
In essence, the reverse mode of automatic differentiation ADreview uses the chain rule of derivative, usually referred to as backpropagation BP1986 in optimization engineering, to compute the gradient information. Formally, in this case
@@
where the $\{T^{(i)},i=1,2,...n\}$ are the intermediate variables which are involved in the process of evaluating $E$, e.g., the temporal local tensors generated in the CTMRG iterations. We assume $n$ related local tensors in Eq. (5) have been generated. In order to compute the gradient $\frac{\partial{E}}{\partial{X}}$, these intermediate tensors generally need to be stored in memory. In the tensor network framework, if the transformations involved in the energy-evaluating contraction algorithms are treated as independent parameters, the automatic differentiation is then equivalent to the backward iteration procedure of the second renormalization group method SRGAD2020 for obtaining the global environment. Therefore, in this sense, this kind of variational tensor network approach tries to solve Eq. (4) by global optimization, and that is one of the reasons why this method can be successfully applied to study a series of strongly correlated systems LW-PTNS2019 , e.g., quantum spin-orbital liquid LY-PRB2023 , altermagnetism LY-PRB2025 , deconfined quantum critical point NingXi-SSM2023 , even with a small bond dimension.
@@ II.2 Review of the NTN method
as illustrated in Fig. 2(a). $T^{b}$, $T^{c}$, $T^{f}$ can be defined similarly corresponding to the $2\times 2$ unit cell in Fig. 1(a). Obviously, each local tensor $T$, and thus the resulting tensor network $\langle\Psi|\Psi\rangle$, has a double-layer structure. That is, if the representation $|\Psi\rangle$ in Eq. (LABEL:Eq:PEPS) has bond dimension $D$, then each index of $T^{a}$ in Eq. (6) and the resulting $\langle\Psi|\Psi\rangle$ have bond dimension $D^{2}$. Therefore, evaluating the energy $E(X)$ requires contracting a tensor network with bond dimension $D^{2}$, which generally leads to a computational cost of about $D^{12}$ NTN2017 when the CTMRG algorithm is employed YJKao-CTMRG ; DNSheng-CTMRG . As explained in Sec. II.1, the memory cost of storing all intermediate tensors generated in this procedure in automatic differentiation is typically $D^{8}$, which would expolding fast with increasing $D$. Therefore, the substantial memory and computational demands impose a stringent constraint on the practically attainable bond dimension $D$ in automatic differentiation-based variational minimization.
@@
In order to solve this problem, the NTN method regards $\langle\Psi|\Psi\rangle$ as a single-layer network with a nesting structure, as illustrated in Fig. 2(c). To be specific, instead of summing over the physical index $\sigma$ and forming a reduced tensor $T^{a}$, as done in Eq. (6), NTN considers the physical indices $\{\sigma\}$ and the link indices equally by regarding $\{\sigma\}$ as special link indices connecting nearby conjugated tensors, e.g, $A^{*}$ in $\langle\Psi|$ and $A$ in $|\Psi\rangle$, as shown in Fig. 2(b). In graphic terms, bra state $\langle\Psi|$ (blue) and ket state $|\Psi\rangle$ (red) lie in the same layer and are connected by $\{\sigma\}$. In order to avoid longer-range links, a virtual tensor $X$, equal to the direct product of two $\delta$ functions, is defined at each crossing point where the virtual links in $\langle\Psi|$ and $|\Psi\rangle$ intersect. Furthermore, to make the resulting network a square lattice, we modify one set of $X$ to $Y$ by combining an $X$ and a $\sigma$. Formally, as shown in Fig. 2(d), $X$ and $Y$ can be defined in the following
@@
Though the idea is simple, the NTN method has a significant advantage that the bond dimension of $\langle\Psi|\Psi\rangle$ is no more than $Dd$, where $d$ is the physical dimension associated with each $\sigma$, instead of $D^{2}$ as in Fig. 2(a). Since the nesting structure is essentially a single-layer tensor network, widely used methods such as CTMRG can be employed to perform the contraction efficiently. For example, suppose the environment dimension used in CTMRG is of the same order as $D^{2}$ required by the reduced tensor network with bond dimension $D^{2}$. In that case, ignoring the scaling of $d$, the contraction of this nested single-layer tensor network has a memory cost of about $D^{6}$ and a computational cost of about $D^{9}$, which are much lower than those in the reduced method ($D^{8}$ and $D^{12}$) NTN2017 .
Bearing this advantage, the NTN method has been used in a series of studies to extend $D$ to larger values on the kagome lattice. For example, $D$ was extended to $25$ in Refs. NTN2017 ; KagomePRL2017 to study the kagome spin liquid, also to $25$ in Ref. YJKao-CTMRG to study the kagome Heisenberg model with DM interactions, and to $15$ in Ref. DNSheng-CTMRG to study the kagome Heisenberg model with chiral interactions.
@@ II.3 The single-layer framework of variational tensor network states
As far as we know, in previous studies, such as Refs. NTN2017 ; KagomePRL2017 ; YJKao-CTMRG ; DNSheng-CTMRG mentioned in Sec. II.2, the NTN method was only utilized in either expectation value calculation or updating the wave function in imaginary-time evolution, but was seldom utilized in the variational calculation of the ground state wave function as discussed in Sec. II.1.
@@
(i). Firstly, we need to prepare an initial PEPS $|\Psi(X_{0})\rangle$ from which the optimization begins. Either an arbitrary state specified by a random $X_{0}$ (this work) or an approximate PEPS obtained by imaginary-time evolution SU ; CU ; FU ; FFU can be accepted.
(ii). Compute the energy $E(X_{0})$ through Eq. (1). where either boundary matrix product state Vidal-TEBD ; Orus-TEBD or CTMRG (this work) can be used to contract the single-layer tensor network $\langle\Psi(X_{0})|\Psi(X_{0})\rangle$ with nested structure explained in Sec. II.2. In this step, one can use automatic differentiation software packages, such as PyTorch Pytorch and Zygote Zygote (this work), to implement the contraction algorithm. Once $E(X_{0})$ is obtained, the gradient $\frac{\partial E(X)}{\partial X}\big|_{X=X_{0}}$ can be evaluated through Eq. (5) efficiently and automatically.
(iii). Using the gradient information obtained in step (ii), we can employ the gradient-based optimization algorithm, such as Adam and L-BFGS OptBook (this work), to get $X_{1}$, the following prediction of the solution of Eq. (4). Then update the parameter by setting $X_{0}=X_{1}$ and obtain a new PEPS $|\Psi(X_{0})\rangle$.
@@
It should be noted that, by converting the reduced the double-layer structure in Fig. 2(a) to the nested single-layer structure in Fig. 2(c), we have made an acceptable sacrifice to reduce the bond dimension $D^{2}$ to $Dd$: the size of the unit cell is doubled from $2\times 2$ to $4\times 4$. Therefore, to contract the tensor network by CTMRG, e.g., one has to employ the multisite version of the algorithm adapted to the nested structure. The specific version of the CTMRG algorithm used in this work, for a $4\times 4$ unit cell, is illustrated in Fig. 3, which is a simple variant of the algorithm proposed in Ref. Corboz-MulCTMRG .
Figure 3: A sketch of the CTMRG algorithm used in this work, for a tensor network with a $4\times 4$ unit cell. The superscripts $i$ and $j$ should be taken modulo 4. (a) The environment of the local tensor located at the $i$-th row and $j$-th column in the unit cell. (b) The recursive relation between the environment of $T^{(i,j)}$ and $T^{(i+1,j)}$, in a left-move. (c) The $4\times 4$ cluster used to determine the transformation matrices used in (b). Here $P^{(i,j)}$ and $Q^{(i+1,j)}$ will be determined, as expressed in Eqs. (8-9), and be inserted in the bonds denoted by $a_{5}a_{6}$. The resulting framework constitutes an optimal approximation of the bond density matrix $\rho$, which has subscripts $(a_{1}a_{2},a_{3}a_{4})$.
Specifically, suppose the distinct tensors in a unit cell are denoted as $\{T^{(i,j)},i,j=0,1,2,3\}$, which are all rank-4 tensors with bond dimension $D$ Explain . In the multisite CTMRG method, the environment of the tensor $T^{(i,j)}$ located at the $i$-th row and $j$-th column in the $4\times 4$ unit cell is represented as $\{C^{(i,j;\alpha)},E^{(i,j;\beta)}\}$, where $\alpha$ and $\beta$ run from 1 to 4 and corresponds to the four corner and four edge tensors for each $T^{(i,j)}$, as illustrated in Fig. 3(a). Therefore, in this case we have 64 distinct corners and 64 edges in total, stored independently for each $T^{(i,j)}$ but updated in a correlated manner. For example, in Fig. 3(b) we explicitly show the recursive relations between the environments of $T^{(i,j)}$ and $T^{(i,j+1)}$, which are used in a left-move step in the CTMRG iterations.
The core step of CTMRG is to find the transformation matrices $P^{(i,j)}$ and $Q^{(i,j)}$, which are of dimension $D\chi$ by $\chi$. Here $\chi$ is the so-called environment dimension, which is the dimension of the bonds belonging only to the environment tensors. Following the bond density matrix approach developed in Refs. HOTRG ; Corboz-MulCTMRG , we always use a $2\times 2$ system surrounded by its environment to determine $P$ and $Q$, as illustrated in Fig. 3(c), where the $4\times 4$ block is intentionally splited into two parts from right. Mathematically, $P^{(i,j)}$ and $Q^{(i+1,j)}$ are constructed in the following
@@
In a single step of left move, for a $4\times 4$ unit cell, we need to determine four pairs of transformation matrices by performing Eqs. (8-9) for each $i$ independently. And after that, one can use these transformations to update the related three environment tensors for each $i$ by performing the contractions illustrated in Fig. 3(b). This finishes a single step of the left move. We need to perform such a step for each $j$, which means we have to perform four such steps in a single left move. Right, upward, and downward moves can be performed one by one similarly, and this completes a single CTMRG iteration. This iteration can be repeated once again, until all the spectra $\tilde{\Lambda}$s converge to some accepted accuracy.
From Fig. 3 and the analysis described above, it is easily shown that if the environment dimension $\chi$ is set to the order of $D^{2}$ empirically NTN2017 , and no symmetry or other advanced techniques like partial matrix decompositions are employed, then the above CTMRG algorithm utilized to contract a nested tensor network with bond dimension $D$ has a memory cost about $D^{6}$ and a computational cost about $D^{9}$ YJKao-CTMRG ; DNSheng-CTMRG . This is more economical than the conventional reduced tensor network method as mentioned in Sec. II.2 and discussed in detail in Ref. NTN2017 . And this is the reason why we can extend our variational calculation equipped with automatic differentiation to $D=9$ in this work.

### v2

# Single-layer framework of variational tensor network states

@@ Abstract
We propose a single-layer tensor network framework for the variational determination of ground states in two-dimensional quantum lattice models. By combining the nested tensor network method [Phys. Rev. B 96, 045128 (2017)] with the automatic differentiation technique, our approach can reduce the computational cost by three orders of magnitude in bond dimension, and therefore enables highly efficient variational ground-state calculations. We demonstrate the capability of this framework through two quantum spin models: the antiferromagnetic Heisenberg model on a square lattice and the frustrated Shastry-Sutherland model. Even without GPU acceleration or symmetry implementation, we have achieved a bond dimension of nine and obtained accurate ground-state energy and consistent order parameters compared to prior studies. In particular, we confirm the existence of an intermediate empty-plaquette valence bond solid ground state in the Shastry-Sutherland model. We have further discussed the convergence of the algorithm and its potential improvements. Our work provides a promising route for large-scale tensor network calculations of two-dimensional quantum systems.
@@ I Introduction
Two-dimensional (2D) strongly correlated systems, such as interacting fermionic systems and frustrated quantum magnets, pose significant challenges for theoretical studies. Consequently, developing efficient and accurate numerical methods for these models constitutes a crucial branch of theoretical condensed matter physics. In recent decades, tensor network methods [1, 2] have attracted increasing attention. By combining the tensor-network ansatz [3, 4] stemming from the quantum entanglement, the contraction strategy originating from renormalization group [6, 5], and modern optimization techniques developed in artificial intelligence [7], tensor network methods have been successfully applied to diverse phenomena, including quantum magnetism [8, 9, 10, 11], superconductivity [12, 13, 14], topological order [15, 16], and quantum field theory [17, 18].
For studying ground states of 2D quantum systems using tensor networks, three strategies are commonly adopted. The first maps the 2D quantum model to a three-dimensional classical one [19], converting the quantum expectation value calculation into a statistical physics problem [20]. The second employs imaginary-time evolution with wave function update schemes [21, 22, 23, 24] to obtain the ground state $|\Psi_{g}\rangle$ in the tensor-network representation, and then evaluate the expectation value $\langle\hat{O}\rangle$ of the operator $\hat{O}$ by contracting the related tensor networks. The third one is the straightforward variational minimization: Starting from an arbitrary tensor network state $|\Psi\rangle$, the ground state of the system $|\Psi_{g}\rangle$ and its energy $E_{g}$ are obtained by minimizing the energy estimator
@@
with respect to tensor parameters [3]. Recently, this approach has become increasingly popular because of the advent of the automatic differentiation technique [25], which, as a mature and widely applied algorithm [26], can efficiently calculate the gradient $\frac{dE}{d|\Psi\rangle}$ [27] and easily integrate with gradient-based optimization method [7] via various open-source packages [28, 29].
@@
A central challenge in all three approaches is the contraction of tensor networks. Unfortunately, contracting a general tensor network in more than one dimension is a #P\#P-complete hard problem [30] that cannot be performed exactly for an infinite system even with translational invariance. To this end, a series of approximation algorithms based on the renormalization group idea have been developed [5, 1], among which the corner transfer-matrix renormalization group (CTMRG) [31, 32, 14, 33] has been widely used. In two dimensions, the cost of the CTMRG algorithm scales as $D^{10}$ in memory and $D^{12}$ in computation with the bond dimension $D$ when no symmetry or partial matrix decomposition is employed. The cost is extremely high for variational calculations, which typically require hundreds of rounds of such contractions. The problem becomes even more serious for minimization with the automatic differentiation since it needs huge memory to store intermediate data for the gradient ($\frac{dE}{d|\Psi\rangle}$) calculation. This issue pose severe limitation for the bond dimension $D$ in variational minimization employing automatic differentiation.
@@
To overcome this limitation, here we propose a single-layer framework of variational tensor network state calculation that incorporates the nested tensor network (NTN) method [34]. By reformulating the double-layer structures of numerator and denominator in Eq. (1) into nested single-layer structures, the NTN method reduces the memory and computation cost of tensor network contraction by two and three orders of $D$, respectively. This allows us to perform variational minimization with significantly larger bond dimension. Our proposal works as follows: For any given state $|\Psi\rangle$, we utilize the NTN to calculate the energy defined in Eq. (1), then calculate the gradient $\frac{dE}{d|\Psi\rangle}$, and finally update the wave function $|\Psi\rangle$ using the gradient information as usual [27]. This procedure can be iterated many times until the calculated energy converges.
@@
We validate this method on two benchmark models defined on infinite 2D lattices: the $S=1/2$ square-lattice antiferromagnetic Heisenberg model and the frustrated Shastry-Sutherland model. The former has a long-range antiferromagnetically ordered ground state [35], while the latter hosts a series of ground states when tuning the strength of spin frustration $J/J^{\prime}$ [36, 10]. In both cases, with the help of NTN, we can efficiently reach $D=9$ even without any GPU utilization or symmetry implementation, which is difficult in previous variational calculations [27, 10, 11]. The obtained ground-state energy and order parameters are consistent with previous calculations, but with better convergence. In particular, we provide clear evidence for an empty-plaquette valence bond solid (VBS) phase in the ground-state phase diagram of the Shastry-Sutherland model.
@@
The rest of the paper is organized as follows. In Sec. II we first briefly review the automatic differentiation-assisted variational calculation of tensor network state and the NTN method, then introduce our single-layer framework. In Sec. III, we present the numerical results of the Heisenberg model and the Shastry-Sutherland model in detail. Finally, we summarize and conclude in Sec. IV.
@@ II Method
In this section, we first review the basics of automatic differentiation-assisted variational calculation of tensor network representation of the ground state [27], the underlying double-layer structure in Eq. (1), and the NTN method [34], then introduce the single-layer framework by combining the two parts together.
@@ II.1 Variational calculation assisted by automatic differentiation
In this work, we follow the variational minimization approach outlined in Sec. I. Starting from an initial $|\Psi(X_{0})\rangle$, we can use well-established techniques, such as the CTMRG [31, 32, 14, 33], to compute the energy $E(X)$. In this procedure, when the automatic differentiation technique is used. Then, as long as $E$ is obtained successfully, we can calculate the gradient $\frac{\partial{E(X)}}{\partial{X}}\big|_{X=X_{0}}$ conveniently by simply calling some open source packages, such as PyTorch [28] and Zygote [29], without any extra operation. After the gradient information is obtained, we can update $X$ using classic optimization algorithms, such as L-BFGS [7], to get a better trial $X_{1}$ and the corresponding wave function $|\Psi(X_{1})\rangle$. Similary, we can evaluate $E(X_{1})$, calculate $\frac{\partial{E(X)}}{\partial{X}}\big|_{X=X_{1}}$, and then obtain a better $X_{2}$ and $|\Psi(X_{2})\rangle$. This iteration can be repeated as many as we want until the gradient $\frac{\partial{E(X)}}{\partial{X}}$ is sufficiently small or the energy $E$ converges within some tolerance. The converged $X^{*}$ yields an approximate solution of Eq. (4) and thereby provides an approximate representation of $|\Psi_{g}\rangle$.
In essence, the reverse mode of automatic differentiation [25] uses the chain rule of derivative, usually referred to as backpropagation [37] in optimization engineering, to compute the gradient information. Formally, in this case
@@
where the $\{T^{(i)},i=1,2,...n\}$ are the intermediate variables that are involved in the process of evaluating $E$, e.g., the temporary local tensors generated in the CTMRG iterations. We assume that $n$ related local tensors in Eq. (5) have been generated. In order to compute the gradient $\frac{\partial{E}}{\partial{X}}$, these intermediate tensors generally need to be stored in memory. In the tensor network framework, if the transformations involved in the energy-evaluation contraction algorithms are treated as independent parameters, automatic differentiation is then equivalent to the backward iteration procedure of the second renormalization group method [38] for obtaining the global environment. Therefore, in this sense, this kind of variational tensor network approach tries to solve Eq. (4) by global optimization, and that is one of the reasons why this method can be successfully applied to study a series of strongly correlated systems [27], e.g., quantum spin-orbital liquid [39], altermagnetism [11], deconfined quantum critical point [10], even with a small bond dimension.
@@ II.2 Review of the NTN method
as illustrated in Fig. 2(a). $T^{b}$, $T^{c}$, $T^{f}$ can be defined similarly corresponding to the $2\times 2$ unit cell in Fig. 1(a). Obviously, each local tensor $T$, and thus the resulting tensor network $\langle\Psi|\Psi\rangle$, has a double-layer structure. That is, if the representation $|\Psi\rangle$ in Eq. (LABEL:Eq:PEPS) has bond dimension $D$, then each index of $T^{a}$ in Eq. (6) and the resulting $\langle\Psi|\Psi\rangle$ have bond dimension $D^{2}$. Therefore, evaluating the energy $E(X)$ requires contracting a tensor network with bond dimension $D^{2}$, which generally leads to a computational cost of about $D^{12}$ [34] when the CTMRG algorithm is employed [40, 41]. As explained in Sec. II.1, the memory cost of storing all intermediate tensors generated in this procedure in automatic differentiation is typically $D^{8}$, which would expand fast with increasing $D$. Therefore, the substantial memory and computational demands impose a stringent constraint on the practically attainable bond dimension $D$ in automatic differentiation-based variational minimization.
@@
In order to solve this problem, the NTN method regards $\langle\Psi|\Psi\rangle$ as a single-layer network with a nesting structure, as illustrated in Fig. 2(c). To be specific, instead of summing over the physical index $\sigma$ and forming a reduced tensor $T^{a}$, as done in Eq. (6), NTN considers the physical indices $\{\sigma\}$ and the link indices equally by regarding $\{\sigma\}$ as special link indices connecting nearby conjugated tensors, e.g, $A^{*}$ in $\langle\Psi|$ and $A$ in $|\Psi\rangle$, as shown in Fig. 2(b). In graphic terms, the bra state $\langle\Psi|$ (blue) and the ket state $|\Psi\rangle$ (red) lie in the same layer and are connected by $\{\sigma\}$. In order to avoid longer-range links, a virtual tensor $X$, equal to the direct product of two $\delta$ functions, is defined at each crossing point where the virtual links in $\langle\Psi|$ and $|\Psi\rangle$ intersect. Furthermore, to make the resulting network a square lattice, we modify one set of $X$ to $Y$ by combining an $X$ and a $\sigma$. Formally, as shown in Fig. 2(d), $X$ and $Y$ can be defined as follows
@@
Though the idea is simple, the NTN method has a significant advantage that the bond dimension of $\langle\Psi|\Psi\rangle$ is no more than $Dd$, where $d$ is the physical dimension associated with each $\sigma$, instead of $D^{2}$ as in Fig. 2(a). Since the nesting structure is essentially a single-layer tensor network, widely used methods such as CTMRG can be employed to perform the contraction efficiently. For example, suppose the environment dimension used in CTMRG is of the same order as $D^{2}$ required by the reduced tensor network with bond dimension $D^{2}$. In that case, ignoring the scaling of $d$, the contraction of this nested single-layer tensor network has a memory cost of about $D^{6}$ and a computational cost of about $D^{9}$, which are much lower than those in the reduced method ($D^{8}$ and $D^{12}$) [34].
Bearing this advantage, the NTN method has been used in a series of studies to extend $D$ to larger values on the kagome lattice. For example, $D$ was extended to $25$ in Refs. [34, 8] to study the kagome spin liquid, also to $25$ in Ref. [40] to study the kagome Heisenberg model with DM interactions, and to $15$ in Ref. [41] to study the kagome Heisenberg model with chiral interactions.
@@ II.3 The single-layer framework of variational tensor network states
As far as we know, in previous studies, such as Refs. [34, 8, 40, 41] mentioned in Sec. II.2, the NTN method was only utilized in either expectation value calculation or updating the wave function in imaginary-time evolution, but was seldom utilized in the variational calculation of the ground state wave function as discussed in Sec. II.1.
@@
(i). First, we need to prepare an initial PEPS $|\Psi(X_{0})\rangle$ from which the optimization begins. Either an arbitrary state specified by a random $X_{0}$ (this work) or an approximate PEPS obtained by imaginary-time evolution [21, 22, 23, 24] can be accepted.
(ii). Compute the energy $E(X_{0})$ through Eq. (1). where either boundary matrix product state [42, 43] or CTMRG (this work) can be used to contract the single-layer tensor network $\langle\Psi(X_{0})|\Psi(X_{0})\rangle$ with nested structure explained in Sec. II.2. In this step, one can use automatic differentiation software packages, such as PyTorch [28] and Zygote [29] (this work), to implement the contraction algorithm. Once $E(X_{0})$ is obtained, the gradient $\frac{\partial E(X)}{\partial X}\big|_{X=X_{0}}$ can be evaluated through Eq. (5) efficiently and automatically.
(iii). Using the gradient information obtained in step (ii), we can employ gradient-based optimization algorithm, such as Adam and L-BFGS [7] (this work), to get $X_{1}$, the updated estimate of the solution of Eq. (4). Then update the parameter by setting $X_{0}=X_{1}$ and obtain a new PEPS $|\Psi(X_{0})\rangle$.
@@
It should be noted that, by converting the reduced double-layer structure in Fig. 2(a) to the nested single-layer structure in Fig. 2(c), we have made an acceptable sacrifice to reduce the bond dimension $D^{2}$ to $Dd$: the size of the unit cell is doubled from $2\times 2$ to $4\times 4$. Therefore, to contract the tensor network by CTMRG, e.g., one has to employ the multisite version of the algorithm adapted to the nested structure. The specific version of the CTMRG algorithm used in this work, for a $4\times 4$ unit cell, is illustrated in Fig. 3, which is a simple variant of the algorithm proposed in Ref. [14].
Figure 3: A sketch of the CTMRG algorithm used in this work, for a tensor network with a $4\times 4$ unit cell. The superscripts $i$ and $j$ should be taken modulo 4. (a) The environment of the local tensor located at the $i$ th row and $j$ th column in the unit cell. (b) The recursive relation between the environment of $T^{(i,j)}$ and $T^{(i+1,j)}$, in a left-move. (c) The $4\times 4$ cluster used to determine the transformation matrices used in (b). Here $P^{(i,j)}$ and $Q^{(i+1,j)}$ will be determined, as expressed in Eqs. (8-9), and be inserted in the bonds denoted by $a_{5}a_{6}$. The resulting framework constitutes an optimal approximation of the bond density matrix $\rho$, which has subscripts $(a_{1}a_{2},a_{3}a_{4})$.
Specifically, suppose the distinct tensors in a unit cell are denoted as $\{T^{(i,j)},i,j=0,1,2,3\}$, which are all rank-4 tensors with bond dimension $D$ [44]. In the multisite CTMRG method, the environment of the tensor $T^{(i,j)}$ located at the $i$ th row and $j$ th column in the $4\times 4$ unit cell is represented as $\{C^{(i,j;\alpha)},E^{(i,j;\beta)}\}$, where $\alpha$ and $\beta$ run from 1 to 4 and correspond to the four corner and four edge tensors for each $T^{(i,j)}$, as illustrated in Fig. 3(a). Therefore, in this case we have 64 distinct corners and 64 edges in total, stored independently for each $T^{(i,j)}$ but updated in a correlated manner. For example, in Fig. 3(b) we explicitly show the recursive relations between the environments of $T^{(i,j)}$ and $T^{(i,j+1)}$, which are used in a left-move step in the CTMRG iterations.
The core step of CTMRG is to find the transformation matrices $P^{(i,j)}$ and $Q^{(i,j)}$, which are of dimension $D\chi$ by $\chi$. Here $\chi$ is the so-called environment dimension, which is the dimension of the bonds belonging only to the environment tensors. Following the bond density matrix approach developed in Refs. [20, 14], we always use a $2\times 2$ system surrounded by its environment to determine $P$ and $Q$, as illustrated in Fig. 3(c), where the $4\times 4$ block is intentionally split into two parts from right. Mathematically, $P^{(i,j)}$ and $Q^{(i+1,j)}$ are constructed in the following
@@
In a single step of left move, for a $4\times 4$ unit cell, we need to determine four pairs of transformation matrices by performing Eqs. (8-9) for each $i$ independently. And after that, one can use these transformations to update the related three environment tensors for each $i$ by performing the contractions illustrated in Fig. 3(b). This finishes a single step of the left move. We need to perform such a step for each $j$, which means we have to perform four such steps in a single left move. Right, upward, and downward moves can be performed one by one similarly, and this completes a single CTMRG iteration. This iteration can be repeated, until all the spectra $\tilde{\Lambda}$s converge to some accepted accuracy.
From Fig. 3 and the analysis described above, it is easily shown that if the environment dimension $\chi$ is set to the order of $D^{2}$ empirically [34], and no symmetry or other advanced techniques like partial matrix decompositions are employed, then the above CTMRG algorithm utilized to contract a nested tensor network with bond dimension $D$ has a memory cost about $D^{6}$ and a computational cost about $D^{9}$ [40, 41]. This is more economical than the conventional reduced tensor network method as mentioned in Sec. II.2 and discussed in detail in Ref. [34]. And this is the reason why we can extend our variational calculation equipped with automatic differentiation to $D=9$ in this work. As also discussed in Ref. [34], as long as $\chi$ is sufficiently large, the CTMRG contraction of a nested tensor network (i.e., single-layer) should produce the same result as that of the corresponding reduced tensor network (i.e., double-layer). We verify this by studying the VBS phase of the Shastry-Sutherland model, and the result is shown in the Appendix.